# Gradient Alignment Improves Test-Time Adaptation for Medical Image Segmentation

### v3

@@ Abstract
Although recent years have witnessed significant advancements in medical image segmentation, the pervasive issue of domain shift among medical images from diverse centres hinders the effective deployment of pre-trained models. Many Test-time Adaptation (TTA) methods have been proposed to address this issue by fine-tuning pre-trained models with test data during inference. These methods, however, often suffer from less-satisfactory optimization due to suboptimal optimization direction (dictated by the gradient) and fixed step-size (predicated on the learning rate). In this paper, we propose the Gradient alignment-based Test-time adaptation (GraTa) method to improve both the gradient direction and learning rate in the optimization procedure. Unlike conventional TTA methods, which primarily optimize the pseudo gradient derived from a self-supervised objective, our method incorporates an auxiliary gradient with the pseudo one to facilitate gradient alignment. Such gradient alignment enables the model to excavate the similarities between different gradients and correct the gradient direction to approximate the empirical gradient related to the current segmentation task. Additionally, we design a dynamic learning rate based on the cosine similarity between the pseudo and auxiliary gradients, thereby empowering the adaptive fine-tuning of pre-trained models on diverse test data. Extensive experiments establish the effectiveness of the proposed gradient alignment and dynamic learning rate and substantiate the superiority of our GraTa method over other state-of-the-art TTA methods on a benchmark medical image segmentation task. The code and weights of pre-trained source models are available at https://github.com/Chen-Ziyang/GraTa.
@@ Introduction
Table 1: Performance of our GraTa, ‘No Adapt’ baseline, and eight competing TTA methods. The best results are highlighted in bold.
@@
We trained a ResUNet-34 (He et al. 2016) backbone following (Hu, Liao, and Xia 2022) as the baseline individually on each domain (source domain) and subsequently tested it on each remaining domain (target domain), computing the mean metrics to evaluate all the methods across diverse scenarios ($4\times 5$ in total). For a fair comparison, we conducted single-iteration adaptation for each batch of test data using a batch size of 1 across all experiments following (Yang et al. 2022). To deploy our GraTa, we utilized the Adam (Kingma and Ba 2014) optimizer to train the affine-transformation parameters within pre-trained models, and the BN statistics were recollected from test data following (Wang et al. 2021; Yang et al. 2022). The scaling factor $\beta$ is set to 0.0001 empirically.
Figure 3: Qualitative results of our GraTa and eight TTA methods using Domain D as the source domain and the remaining domains as target domains. We displayed the results of two simple samples (rows 1-2) and two hard samples (rows 3-4). The ground-truth bounding boxes of OD and OC were overlaid on each segmentation result to highlight potential over- or under-segmentation. Best viewed in color.
@@ Results
We compare our GraTa with the ‘No Adapt’ baseline (testing without adaptation), and eight competing TTA methods, including three methods based on refining the BN statistics (DUA (Mirza et al. 2022), DIGA (Wang et al. 2023b), and MedBN (Park et al. 2024)), and five methods based on fine-tuning the pre-trained models (TENT (Wang et al. 2021), DLTTA (Yang et al. 2022), DomainAdaptor (Zhang et al. 2023), SAR (Niu et al. 2023), and DeTTA (Wen et al. 2024)). Specifically, we re-implemented all competing methods using the same baseline as our GraTa and integrated DLTTA with the entropy loss proposed in TENT (Yang et al. 2022). The result of each domain is the average metric calculated by treating the current domain as the source domain and employing the pre-trained model on the remaining domains.
@@
Comparing to other TTA methods. The results of our GraTa, the “No Adapt” baseline, and eight competing TTA methods are detailed in Table 1. The BN-based methods exhibit stable but suboptimal enhancements across all domains due to the absence of updating the model parameters. TENT exhibits limited performance, while other variants (i.e., DLTTA, DomainAdaptor, and SAR) enhance the performance significantly. It can be observed that our GraTa achieves the best performance across most scenarios and improves the baseline by $5.87\%$, which demonstrates the effectiveness and superiority of our GraTa. Specifically, our GraTa boosts the performance in Domain D by $12.77\%$, while other methods exhibit marginal improvement, which means our approach extends the upper limit of TTA in challenging scenarios.
@@
Ablation study. To evaluate the contributions of our optimization objective and dynamic learning rate, we conducted a series of ablation experiments, as shown in Table 2. It shows that (1) utilizing $\mathcal{L}_{con}(\theta;\mathcal{X}^{t}_{i})$ only obtains limited performance similar to TENT; (2) introducing $\mathcal{L}_{con}(\theta-\nabla\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ and $\eta$ enhances the performance by $2.14\%$ and $2.81\%$ respectively; (3) the best performance is achieved when both are utilized jointly (i.e., our GraTa).
Qualitative analysis. We selected the model trained on Domain D (source) as a case study and visualized four test images from four scenarios, denoted as “D $\rightarrow$ target domain,” along with their corresponding segmentation results produced by our GraTa method, eight competing methods, and the ground truth, as shown in Figure 3. The regions of the OD and OC are highlighted in green and blue, respectively. For comparative purposes, we drew the bounding boxes of the OD and OC on the ground truth and overlaid them on each segmentation result. Compared to other competing methods, our GraTa method demonstrates reduced over-segmentation and under-segmentation across all samples, indicating the effectiveness of the improved optimization procedure.
@@
Exchange the Location of $\mathcal{L}_{con}$ and $\mathcal{L}_{ent}$. It is particularly critical to determine which objective function is the optimal one to produce the pseudo gradient. We exchanged the location of $\mathcal{L}_{con}$ and $\mathcal{L}_{ent}$ in our objective and repeated the experiments for evaluation. The results are listed in Table 3. Compared to TENT which optimizes $\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i})$ only, the variant of our GraTa using $\mathcal{L}_{ent}(\theta-\nabla\mathcal{L}_{con}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ as the optimization objective and $\eta$ as the learning rate also boosts the performance. Our GraTa achieves the best performance, which demonstrates that $\mathcal{L}_{con}$ is more suitable as the objective function to produce the pseudo gradient due to its superior supervision.
Obtain $\eta$ using other activation functions. To ascertain the optimal activation function for mapping the cosine similarity to obtain $\eta$, we conducted experiments utilizing various common activation functions and listed the results in Table 4. The results reveal that (1) our designed function achieves superior performance to others by amplifying the distinctions among learning rates generated by diverse cosine similarity scores; (2) ReLU exhibits the worst average performance which underscores the necessity of assigning appropriate learning rates for fine-tuning even if distinct gradients diverge largely.
@@
More combinations of objective functions. For further discussion, we introduced four additional loss functions, i.e., the reconstruction loss, the rotation prediction loss, the super-resolution loss, and the denoising loss. We utilized the cross-entropy loss to calculate the rotation prediction loss and the mean-square-error loss to calculate others. To construct the super-resolution loss, we downsampled the input image by a factor of four. For the denoising loss, we added the Gaussian noise sampled from a standard normal distribution to the image. We evaluated all possible combinations of them to determine the optimal one and displayed the results in Table 5. It reveals that (1) using $\mathcal{L}_{con}$ as the pseudo objective function achieves superior performance over others; (2) using $\mathcal{L}_{rec}$, $\mathcal{L}_{rot}$, $\mathcal{L}_{sup}$, or $\mathcal{L}_{den}$ as the pseudo objective function results in similar but suboptimal performance since they all require an additional branch to construct the loss, making it challenging provide appropriate gradients during inference; (3) using $\mathcal{L}_{rec}$, $\mathcal{L}_{rot}$, $\mathcal{L}_{sup}$, or $\mathcal{L}_{den}$ as the auxiliary objective function can enhance the performance of $\mathcal{L}_{con}$ and $\mathcal{L}_{ent}$, demonstrating the effectiveness of our GraTa.
@@ Conclusion
In this paper, we proposed the GraTa, a TTA method to focus on optimizing the optimization direction and step-size during inference by gradient alignment and utilizing a dynamic learning rate, respectively. Specifically, we introduced an auxiliary gradient to the pseudo one to perform alignment to excavate the similarities between distinct gradients and further approximate the empirical gradient. The dynamic learning rate is constructed based on the cosine similarity between the pseudo and auxiliary gradients, encouraging pre-trained models to enhance learning as these gradients become closely aligned. Extensive experiment results on diverse scenarios demonstrated the superiority of our GraTa and the effectiveness of each special design. In our future work, we intend to explore the alignment of other auxiliary gradients to further enhance performance, as well as investigate other deployment scenarios, such as continual test-time adaptation (Chen et al. 2024). We hope that our efforts can promote research on gradient optimization for test-time adaptation.

### v4

@@ Abstract
Although recent years have witnessed significant advancements in medical image segmentation, the pervasive issue of domain shift among medical images from diverse centres hinders the effective deployment of pre-trained models. Many Test-time Adaptation (TTA) methods have been proposed to address this issue by fine-tuning pre-trained models with test data during inference. These methods, however, often suffer from less-satisfactory optimization due to suboptimal optimization direction (dictated by the gradient) and fixed step-size (predicated on the learning rate). In this paper, we propose the Gradient alignment-based Test-time adaptation (GraTa) method to improve both the gradient direction and learning rate in the optimization procedure. Unlike conventional TTA methods, which primarily optimize the pseudo gradient derived from a self-supervised objective, our method incorporates an auxiliary gradient with the pseudo one to facilitate gradient alignment. Such gradient alignment enables the model to excavate the similarities between different gradients and correct the gradient direction to approximate the empirical gradient related to the current segmentation task. Additionally, we design a dynamic learning rate based on the cosine similarity between the pseudo and auxiliary gradients, thereby empowering the adaptive fine-tuning of pre-trained models on diverse test data. Extensive experiments establish the effectiveness of the proposed gradient alignment and dynamic learning rate and substantiate the superiority of our GraTa method over other state-of-the-art TTA methods on a benchmark medical image segmentation task.
@@ Introduction
Table 1: Performance of our GraTa, ‘No Adapt’ baseline, and eight TTA methods. The best results are highlighted in bold.
@@
We trained a ResUNet-34 (He et al. 2016) backbone following (Hu, Liao, and Xia 2022) as the baseline individually on each domain (source domain) and subsequently tested it on each remaining domain (target domain), computing the mean metrics to evaluate all the methods across diverse scenarios ($4\times 5$ in total). For a fair comparison, we conducted single-iteration adaptation for each batch of test data using a batch size of 1 across all experiments following (Yang et al. 2022). To deploy our GraTa, we utilized the Adam (Kingma and Ba 2014) optimizer to train the affine-transformation parameters within pre-trained models, and the BN statistics were recollected from test data following (Wang et al. 2021). The scaling factor $\beta$ is set to 0.0001 empirically.
Figure 3: Qualitative results of our GraTa and eight TTA methods using Domain D as the source domain and remaining domains as target domains. We displayed the results of two simple samples (rows 1-2) and two hard samples (rows 3-4). The ground-truth bounding boxes of OD and OC were overlaid on each segmentation result to highlight potential over- or under-segmentation. Best viewed in color.
@@ Results
We compare our GraTa with the ‘No Adapt’ baseline (testing without adaptation), and eight competing TTA methods, including three methods based on refining the BN statistics (DUA (Mirza et al. 2022), DIGA (Wang et al. 2023b), and MedBN (Park et al. 2024)), and five methods based on fine-tuning the pre-trained models (TENT (Wang et al. 2021), DLTTA (Yang et al. 2022), DomainAdaptor (Zhang et al. 2023), SAR (Niu et al. 2023), and DeTTA (Wen et al. 2024)). Specifically, we re-implemented all competing methods using the same baseline as our GraTa and integrated DLTTA with the entropy loss proposed in TENT (Yang et al. 2022). The result of each domain is the average metric calculated by treating the current domain as the source domain and employing the pre-trained model on remaining domains.
@@
Comparing to other TTA methods. The results of our GraTa, the “No Adapt” baseline, and eight competing TTA methods are detailed in Table 1. The BN-based methods exhibit stable but suboptimal enhancements across all domains due to the absence of model updating. TENT exhibits limited performance, while other variants (i.e., DLTTA, DomainAdaptor, and SAR) enhance the performance significantly. It can be observed that our GraTa achieves the best performance across most scenarios and improves the baseline by $5.87\%$, which demonstrates the effectiveness and superiority of our GraTa. Specifically, our GraTa boosts the performance in Domain D by $12.77\%$, while other methods exhibit marginal improvement, which means our approach extends the upper limit of TTA in challenging scenarios.
@@
Ablation study. To evaluate the contributions of our optimization objective and dynamic learning rate, we conducted a series of ablation experiments, as shown in Table 2. It shows that (1) using $\mathcal{L}_{con}(\theta;\mathcal{X}^{t}_{i})$ only obtains limited performance similar to TENT; (2) introducing $\mathcal{L}_{con}(\theta-\nabla\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ and $\eta$ enhances the performance by $2.14\%$ and $2.81\%$, respectively; (3) the best performance is achieved when both are utilized jointly (i.e., our GraTa).
Qualitative analysis. We selected the model trained on Domain D (source) as a case study and visualized four test images from four scenarios, denoted as “D $\rightarrow$ target domain,” along with their corresponding segmentation results produced by our GraTa method, eight competing methods, and the ground truth, as shown in Figure 3. The regions of OD and OC are highlighted in green and blue, respectively. For comparative purposes, we drew the bounding boxes of the OD and OC on the ground truth and overlaid them on each segmentation result. Compared to other competing methods, our GraTa demonstrates reduced over-segmentation and under-segmentation across all samples, indicating the effectiveness of the improved optimization procedure.
@@
Exchange the Location of $\mathcal{L}_{con}$ and $\mathcal{L}_{ent}$. It is particularly critical to determine which objective function is the optimal one to produce the pseudo gradient. We exchanged the location of $\mathcal{L}_{con}$ and $\mathcal{L}_{ent}$ in our objective and repeated the experiments for evaluation. The results are listed in Table 3. Compared to TENT which optimizes $\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i})$ only, the variant of our GraTa using $\mathcal{L}_{ent}(\theta-\nabla\mathcal{L}_{con}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ as the optimization objective and $\eta$ as the learning rate also boosts the performance. Our GraTa achieves the best performance, which demonstrates that $\mathcal{L}_{con}$ is more suitable to produce the pseudo gradient due to its superior supervision.
Obtain $\eta$ using other activation functions. To ascertain the optimal activation function for mapping the cosine similarity to obtain $\eta$, we conducted experiments using various activation functions and listed the results in Table 4. It reveals that (1) our designed function achieves superior performance to others by amplifying the distinctions among learning rates generated by diverse cosine similarity scores; (2) ReLU exhibits the worst average performance which underscores the necessity of assigning appropriate learning rates for fine-tuning even if distinct gradients diverge largely.
@@
More combinations of objective functions. For further discussion, we introduced four additional loss functions, i.e., the reconstruction loss, the rotation prediction loss, the super-resolution loss, and the denoising loss. We utilized the cross-entropy loss to calculate the rotation prediction loss and the mean-square-error loss to calculate others. To construct the super-resolution loss, we downsampled the input image by a factor of four. For the denoising loss, we added the Gaussian noise sampled from a standard normal distribution to the image. We evaluated all possible combinations of them to determine the optimal one and displayed the results in Table 5. It reveals that (1) using $\mathcal{L}_{con}$ as the pseudo objective function achieves superior performance over others; (2) using $\mathcal{L}_{rec}$, $\mathcal{L}_{rot}$, $\mathcal{L}_{sup}$, or $\mathcal{L}_{den}$ as the pseudo objective function results in similar but suboptimal performance since they all require an additional branch to construct the loss, making it challenging to provide appropriate gradients during inference; (3) using $\mathcal{L}_{rec}$, $\mathcal{L}_{rot}$, $\mathcal{L}_{sup}$, or $\mathcal{L}_{den}$ as the auxiliary objective function can enhance the performance of $\mathcal{L}_{con}$ and $\mathcal{L}_{ent}$, demonstrating the effectiveness of GraTa.
@@ Conclusion
In this paper, we proposed the GraTa, a TTA method to focus on optimizing the optimization direction and step-size during inference by gradient alignment and utilizing a dynamic learning rate, respectively. Specifically, we introduced an auxiliary gradient to the pseudo one to perform alignment to excavate the similarities between distinct gradients and further approximate the empirical gradient. The dynamic learning rate is constructed based on the cosine similarity between the pseudo and auxiliary gradients, encouraging pre-trained models to enhance learning as these gradients become closely aligned. Extensive experiment results on diverse scenarios demonstrated the superiority of our GraTa and the effectiveness of each special design. In our future work, we intend to explore the alignment of other auxiliary gradients to further enhance performance, as well as investigate other deployment scenarios, such as continual test-time adaptation (Chen et al. 2024). We hope that our efforts can promote research on gradient optimization for TTA.